A wholesaler is offering two different package deals of roses and carnations to florists. One package contains 20 dozen roses and 34 dozen carnations for $504.00. The other package contains 15 dozen roses and 17 dozen carnation for
$327.00. This information can be represents by the system of equations below,where r represents the cost of one dozen roses and c represents the cost of one dozen carnations.
20r+34c=504
15r+17c=327
solve the system of equations to find the cost,in dollar,of a dozen roses

Answer by ionesco(12)

20r+34c=504
15r+17c=327
multiply the second with 2 and substract first from second
so,
20r+34c=504
30r+34c=654
then 10r=150 then a dozen roses will cost 15$.